# A Generalized Information Bottleneck Theory of Deep Learning

### v2

@@ 5.3 Information Dynamics of Transformers Classifying News Headlines
The standard fine-tuning approach shown in row 1 of Figure 6 produces highly variable trajectories that begin near the center of the information plane, indicating that pre-trained BERT representations contain substantial pre-training biases. This prompted us to unlearn, where we train on random labels. The unlearning intervention dramatically alters these dynamics. After 3 epochs of random label training, models consistently start from the bottom-right corner of the information plane, as shown in row 2. From this reset position, both IB and GIB show more coherent learning trajectories during subsequent fine-tuning. This reveals how studying these information-planes can be used for diagnostics and interpretation. It is worth noting that, in the absence of unlearning for seed 2, the model failed to overcome its initial biases, which led to atypical and unstable information dynamics.
@@ Appendix B Proof of Theorem 1
$\mathcal{L}_{\text{GIB}}$ $=I(\mathcal{X};Q(Z,Y))-\frac{1}{2\beta N}\sum_{i=1}^{N}\left(I(\mathcal{X}^{-i};Q(Z,Y))+I(X^{i};Q(Z,Y))\right)$ (9)
$=I(\mathcal{X};Y)-\frac{1}{2\beta N}\sum_{i=1}^{N}\left(I(\mathcal{X}^{-i};Z)+I(X^{i};Z)\right)$ (10)
$\geq I(\mathcal{X};Y)-\frac{1}{2\beta N}\sum_{i=1}^{N}\left(H(Z)+H(Z)\right)\quad\text{(by monotonicity.)}$ (11)
@@
$\geq I(\mathcal{X};Y)-\frac{1}{\beta}I(\mathcal{X};\mathcal{T})\quad\text{(by dpi.)}$ (15)
$\geq I(\mathcal{T};Y)-\frac{1}{\beta}I(\mathcal{X};\mathcal{T})\quad\text{(by dpi.)}$ (16)
@@ Appendix C GIB and Sufficient Statistics
In this section, we prove the GIB discovers sufficient statistics.
@@
Let $(X,Y)\sim P_{X,Y}$ and let $Z\sim P_{Z|X}$ be any (possibly stochastic) encoder. Define
$Q(Z,Y)\;=\;\frac{P_{Z,Y}(Z,Y)}{P_{Z}(Z)P_{Y}(Y)}\;=\;\frac{P_{Z|Y}(Z|Y)}{P_{Z}(Z)}\;=\;\frac{P_{Y|Z}(Y|Z)}{P_{Y}(Y)}.$
For $\beta\in(0,\infty]$, consider
$\mathcal{J}_{\beta}(P_{Z|X})\;:=\;I\!\big(X;Q(Z,Y)\big)-\frac{1}{\beta\,2N}\sum_{i=1}^{N}\!\Big(I\!\big(X^{-i};Q(Z,Y)\big)+I\!\big(X^{i};Q(Z,Y)\big)\Big),$
where $X=(X^{1},\dots,X^{N})$ and $X^{-i}$ omits coordinate $i$. Then, at $\beta=\infty$,
$\sup_{P_{Z|X}}\mathcal{J}_{\infty}\;=\;\sup_{P_{Z|X}}I\!\big(X;Q(Z,Y)\big)\;=\;I(X;Y),$
and the supremum is attained if and only if $Z$ is sufficient for $Y$ given $X$, i.e., $P_{Y|X}=P_{Y|Z}$ almost surely.
@@
At $\beta=\infty$ the penalty vanishes and $\mathcal{J}_{\infty}=I(X;Q)$. The transformation $(X,Y)\mapsto(X,Q)$ is a (possibly randomized) Markov kernel induced by first drawing $Z\sim P_{Z|X}(\cdot|X)$ and then setting $Q=\frac{P_{Y|Z}(Y|Z)}{P_{Y}(Y)}$. By data processing for KL divergence,
$I\!\big(X;Q(Z,Y)\big)\;\leq\;I(X;Y).$
Taking the supremum over $P_{Z|X}$ yields $\sup_{P_{Z|X}}I(X;Q)\leq I(X;Y)$. From this point onward, we adopt the same analytical approach as the IB framework to demonstrate that the GIB identifies sufficient statistics. See Tishby et al. (1999). ∎
@@ Appendix D When GIB Encounters Infinity
Let $\mathcal{X}$ and $Y$ be continuous random variables with continuous probability distributions. Assume perfect training accuracy such that $Q(Z,Y)=Z=Y$ and $H(Z|\mathcal{X})=0$ (deterministic predictor). If no single feature or feature subset can perfectly predict the output, i.e.,
$\frac{1}{N}\sum_{i=1}^{N}\left(I(\mathcal{X}^{-i};Z)+I(X^{i};Z)\right)<\infty,$
then $\mathcal{L}_{\text{GIB}}=\infty$.
@@
Starting from our GIB formulation with $Q(Z,Y)=Z=Y$:
$\mathcal{L}_{\text{GIB}}$ $=I(\mathcal{X};Q(Z,Y))-\frac{1}{2\beta N}\sum_{i=1}^{N}\left(I(\mathcal{X}^{-i};Q(Z,Y))+I(X^{i};Q(Z,Y))\right)$ (18)
$=I(\mathcal{X};Z)-\frac{1}{2\beta N}\sum_{i=1}^{N}\left(I(\mathcal{X}^{-i};Z)+I(X^{i};Z)\right)$ (19)
$=\infty-\frac{1}{2\beta N}\sum_{i=1}^{N}\left(I(\mathcal{X}^{-i};Z)+I(X^{i};Z)\right)$ (20)
$=\infty$ (21)
The third line follows because $I(\mathcal{X};Z)=\infty$ for continuous variables with perfect dependence, while the sum remains finite by assumption. ∎
@@
In contrast to the original IB, where infinities arise as technical artifacts, the infinities in our GIB possess more interpretability. When $\mathcal{L}_{\text{GIB}}=\infty$, we have a case of perfect synergy: the output cannot be predicted without the complete feature set, and removing even a single feature results in loss of predictive power. This represents the ideal synergistic function where all features must interact to determine the output.
Conversely, if any single feature subset can perfectly predict the output (in violation of our assumption), then at least one of the terms $I(\mathcal{X}^{-i};Z)$ or $I(X^{i};Z)$ becomes infinite, making the entire sum infinite. In this case, $\mathcal{L}_{\text{GIB}}=\infty-\infty=0$. Therefore, while both IB and GIB yield infinities in the continuous setting, our formulation admits a meaningful interpretation: an infinite GIB corresponds to perfect synergy, indicating that all features are essential. This contrasts with standard IB, where infinities are just a technical feature of continuous random variables under deterministic functions.
@@ E.6 Adversarial Robustness
Appendix F Alternative MI Estimation Using Loss Comparison
To validate the robustness of our findings, we repeated our experiments using an alternative MI estimation method based on predictive power (Covert et al., 2020). This method estimates MI by training a model to predict one variable from another and measuring the predictive performance. For discrete targets, we use cross-entropy loss to directly recover the MI in bits. For continuous targets, we use mean squared error (MSE) loss, which provides an approximate MI estimate. While this approach is more computationally intensive than histogram binning, it can potentially capture more complex dependencies.
F.1 Activation Function Comparison
Figure 8: Information plane dynamics across activation functions using loss comparison method MI estimation. Standard IB (blue) and GIB (pink) trajectories for networks trained on synthetic tasks S0-S4. MI computed every 10 epochs for first 2000 epochs, then every 250 epochs.
Figure 8 shows information plane dynamics using loss comparison estimation. While compression phases are less clearly defined than with histogram binning, we observe that only our GIB formulation (pink trajectories) exhibits compression phases across the different activation functions. The standard IB (blue trajectories) shows minimal or no compression, consistent with our main findings. The reduced clarity of phases likely stems from the increased variance inherent in the loss comparison method, which requires training auxiliary models at each measurement point.
F.2 Simple Mathematical Functions
Figure 9: Information plane dynamics for simple mathematical functions using loss comparison MI estimation. Due to computational constraints, MI was computed every 10 epochs rather than continuously.
Figure 9 presents results for networks learning arithmetic and polynomial functions. For the addition task, we observe some compression phases in our GIB formulation but not in the standard IB. However, for the remaining functions (multiplication and symmetric polynomials), the information plane trajectories show no clear phase structure for either method. This is likely due to two factors: (1) the use of MSE loss for continuous targets provides only approximate MI estimates, and (2) the reduced sampling frequency (every 10 epochs) may miss rapid transitions. The computational cost of loss comparison necessitated this reduced sampling rate, as each MI estimate requires training an auxiliary model.
F.3 ResNet Information Dynamics
@@
Figure 10: Information plane dynamics for ResNets on CIFAR-10 using loss comparison MI estimation. MI computed at every epoch using the same PCA preprocessing as in main experiments.
Figure 10 reveals an interesting pattern in ResNet training dynamics. The standard IB (blue trajectories) typically shows a clear fitting phase (upward movement) but often lacks a subsequent compression phase. In contrast, our GIB formulation (pink trajectories) exhibits the opposite behavior: less pronounced fitting phases but more consistent compression. This asymmetry suggests that loss comparison estimation may be more sensitive to the synergistic decomposition in our GIB formulation than to the aggregate information flow measured by standard IB.
F.4 BERT Fine-tuning Dynamics
@@
Figure 11: Information plane dynamics for BERT fine-tuning using loss comparison MI estimation. Top row shows standard fine-tuning (Learned), bottom row shows unlearning followed by fine-tuning (Unlearned).
Figure 11 examines BERT fine-tuning with loss comparison estimation. Before the unlearning intervention (top row), both IB and GIB trajectories are difficult to interpret, showing erratic patterns without clear phase structure. After unlearning (bottom row), the dynamics become more structured: both methods consistently show fitting phases across all seeds. However, only the GIB formulation exhibits a compression phase, and this is clearly visible only for seed 1. The increased noise in these measurements compared to histogram binning may reflect the challenge of estimating MI in high-dimensional token spaces using loss comparison.
F.5 Summary of Loss Comparison Results
While loss comparison provides a theoretically principled approach to MI estimation, our results demonstrate several practical limitations. The method produces noisier estimates than histogram binning, particularly for continuous targets and high-dimensional inputs. Despite this increased noise, our GIB formulation was still more likely to exhibit compression phases compared to standard IB, which often showed only fitting phases or no clear dynamics at all. This robustness is particularly notable given the well-documented brittleness of information plane dynamics. As demonstrated by Saxe et al. (2019) and later by Geiger (2022), observing clear compression phases in the standard IB framework depends critically on the choice of activation functions, binning procedures, and even random initialization. The fact that our GIB formulation shows more consistent phase structure across different MI estimation methods may suggest that synergy-based decomposition captures a more fundamental aspect of how neural networks process information during learning.
@@ Appendix G Computational Complexity
The computational requirements of GIB and IB differ significantly in their scaling behavior. If we define our unit of computation as a single MI estimation, GIB requires $2N+1$ calculations: more specifically, one for the prediction term $I(\mathcal{X};Q(Z,Y))$ and $2N$ for the complexity term (computing $I(\mathcal{X}^{-i};Q(Z,Y))$ and $I(X^{i};Q(Z,Y))$ for each feature). Critically, these calculations occur at the input layer where dimensionality is typically highest, for CIFAR-10, this means 3072 features. However, GIB’s computational cost is independent of network depth, since it only considers input-output relationships. In contrast, standard IB requires $2L$ MI calculations for $L$ layers, computing $I(\mathcal{X};\mathcal{T}_{l})$ and $I(\mathcal{T}_{l};Y)$ at each layer. While one might compute IB only for the final layer where dynamics are often most pronounced, this prevents the use of the IB as a tool for understanding learning dynamics throughout the network. Additionally, GIB benefits from a key advantage: we can apply PCA to high-dimensional inputs (as we do for CIFAR-10 in Section 5.2) because features at the input layer share a common representation space (Turk & Pentland, 1991). Conversely, combining representations across layers for IB is less conventional.
@@ Appendix H Effect of PCA Dimensionality on GIB Dynamics
Figure 12 demonstrates the impact of PCA dimensionality on observed GIB dynamics. With only 25 principal components, the compression phases are absent; trajectories show limited leftward movement and often remain clustered. However, as we increase to 50 components (shown in main results) and then to 100 components, the compression phases become increasingly apparent. This progression suggests that capturing synergistic information requires sufficient dimensionality to represent the complex feature interactions present in the original input space.

### v3

@@ 5.3 Information Dynamics of Transformers Classifying News Headlines
The standard fine-tuning approach shown in row 1 of Figure 6 produces highly variable trajectories that begin near the center of the information plane, indicating that pre-trained BERT representations contain substantial pre-training biases. This prompted us to unlearn, where we train on random labels. The unlearning intervention dramatically alters these dynamics. After 3 epochs of random label training, models consistently start from the bottom-right corner of the information plane, as shown in row 2. From this reset position, both IB and GIB show more coherent learning trajectories during subsequent fine-tuning. This reveals how studying these information-planes can be used for diagnostics and interpretation.
@@ Appendix B Proof of Theorem 1
$\mathcal{L}_{\text{GIB}}$ $=I(Z;Y)-\frac{1}{2\beta N}\sum_{i=1}^{N}\left(I(\mathcal{X}^{-i};Q)+I(X^{i};Q)\right)$ (8)
$=H(Y)-\frac{1}{2\beta N}\sum_{i=1}^{N}\left(I(\mathcal{X}^{-i};Z)+I(X^{i};Z)\right)\quad\text{(since $Z=Y$ and $Q=Z=Y$)}$ (9)
$\geq I(\mathcal{X};Y)-\frac{1}{2\beta N}\sum_{i=1}^{N}\left(I(\mathcal{X}^{-i};Z)+I(X^{i};Z)\right)\quad\text{(since $I(\mathcal{X};Y)\leq H(Y)$)}$ (10)
$\geq I(\mathcal{X};Y)-\frac{1}{2\beta N}\sum_{i=1}^{N}\left(H(Z)+H(Z)\right)\quad\text{(by $I(A;B)\leq H(B)$)}$ (11)
@@
$\geq I(\mathcal{X};Y)-\frac{1}{\beta}I(\mathcal{X};\mathcal{T})\quad\text{(by data processing, $\mathcal{X}\to\mathcal{T}\to Z$)}$ (15)
$\geq I(\mathcal{T};Y)-\frac{1}{\beta}I(\mathcal{X};\mathcal{T})\quad\text{(by data processing, $\mathcal{X}\to\mathcal{T}\to Y$).}$ (16)
@@ Appendix C GIB and Sufficient Statistics
In this section, we show that when the first term in the GIB objective is $I(Z;Y)$, the GIB recovers sufficient statistics in the limit $\beta\to\infty$.
@@
For $\beta\in(0,\infty]$, consider the GIB functional
$\mathcal{J}_{\beta}(P_{Z|X}):=I(Z;Y)-\frac{1}{\beta\,2N}\sum_{i=1}^{N}\Big(I(X^{-i};Q)+I(X^{i};Q)\Big).$
Then at $\beta=\infty$,
$\sup_{P_{Z|X}}\mathcal{J}_{\infty}=\sup_{P_{Z|X}}I(Z;Y)\;\leq\;H(Y).$
Moreover, equality holds (i.e., $\sup_{P_{Z|X}}I(Z;Y)=H(Y)$) if and only if $Y$ is a deterministic function of $Z$, equivalently $H(Y\mid Z)=0$. In this case, $Z$ is a sufficient statistic.
@@
At $\beta=\infty$, the penalty term vanishes and therefore $\mathcal{J}_{\infty}=I(Z;Y)$. By the elementary information inequality $I(Z;Y)\leq H(Y)$ for all encoders $P_{Z|X}$, we have
$\sup_{P_{Z|X}}\mathcal{J}_{\infty}=\sup_{P_{Z|X}}I(Z;Y)\leq H(Y).$
If for some encoder we have $H(Y\mid Z)=0$, i.e. $Y$ is a deterministic function of $Z$, then
$I(Z;Y)=H(Y)-H(Y\mid Z)=H(Y),$
and the upper bound is achieved.
Conversely, if $I(Z;Y)=H(Y)$ for some encoder, then
$H(Y\mid Z)=H(Y)-I(Z;Y)=0,$
so $Y$ must be a deterministic function of $Z$.
Thus, at $\beta=\infty$, maximising $\mathcal{J}_{\infty}$ recovers precisely the set of encoders for which $Z$ is a sufficient statistic for $Y$. ∎
@@ Appendix D When GIB Encounters Infinity
Let $\mathcal{X}=(X^{1},\dots,X^{N})$ be a feature vector and let $Y$ be a continuous random variable. Assume that the representation $Z$ achieves perfect prediction of $Y$ in the sense that $I(Z;Y)=\infty.$ Assume further that no individual feature $X^{i}$ nor its complement $\mathcal{X}^{-i}$ alone yields an infinite mutual information with the PMI variable $Q$, in the sense that
$\frac{1}{N}\sum_{i=1}^{N}\Big(I(\mathcal{X}^{-i};Q)+I(X^{i};Q)\Big)<\infty,$
where $Q$ is the point-wise mutual information random variable. Then $\mathcal{L}_{\mathrm{GIB}}=\infty$.
@@
By definition,
$\mathcal{L}_{\mathrm{GIB}}=I(Z;Y)-\frac{1}{2\beta N}\sum_{i=1}^{N}\Big(I(\mathcal{X}^{-i};Q)+I(X^{i};Q)\Big).$
By assumption, $I(Z;Y)=\infty$, while the sum is finite. Hence, in the extended real line,
$\mathcal{L}_{\mathrm{GIB}}=\infty-\text{(finite)}=\infty.$
@@
In contrast to the standard IB, where infinities arise as technical artifacts of deterministic mappings between continuous variables, the infinities in the GIB functional admit a more structural interpretation.
When $\mathcal{L}_{\mathrm{GIB}}=\infty$, we are in a regime of perfect synergy: the output $Y$ can be recovered from the full feature set (via $Z$), yet no individual coordinate or coordinate subset carries enough information to reconstruct the relevant PMI signal. In other words, all features are jointly essential. This corresponds to the ideal synergistic regime.
Conversely, if some feature $X^{i}$ or complement $\mathcal{X}^{-i}$ can alone perfectly determine $Y$, then the corresponding term $I(X^{i};Q)$ or $I(\mathcal{X}^{-i};Q)$ becomes infinite, making the entire penalty infinite. In this case, the GIB takes the form
$\infty-\infty,$
which is indeterminate. This suggests that prediction is not synergistic, as the output can be reconstructed from a sufficiently informative subset of the features.
Thus, while both IB and GIB encounter infinities in the continuous setting, the GIB admits a meaningful interpretation: an infinite value corresponds precisely to perfect synergy, whereas the indeterminate case reflects the absence of synergy.
@@ E.6 Adversarial Robustness
Appendix F Alternative MI Estimation Using GCMI and KDE
To validate the robustness of our findings, we repeated our experiments using two alternative MI estimation methods: Gaussian Copula Mutual Information (GCMI) and Kernel Density Estimation (KDE). These methods offer different trade-offs between computational efficiency and estimation accuracy compared to our primary histogram binning approach.
F.1 Kernel Density Estimation (KDE)
KDE (Parzen, 1962) estimates probability densities using kernel functions centered at each data point, then computes MI from these continuous density estimates. We use Gaussian kernels with bandwidth selected via Scott’s rule (Scott, 1992). While computationally more intensive than GCMI ($\mathcal{O}(n^{2})$ for $n$ samples), KDE provides non-parametric estimates that can capture arbitrary distribution shapes without assuming specific parametric forms. This flexibility makes KDE particularly suitable for complex, multi-modal distributions that might arise in neural network representations.
F.1.1 Activation Function Comparison
Figure 8: Information plane dynamics across activation functions using KDE estimation. The non-parametric nature of KDE reveals fine-grained dynamics in the information plane trajectories.
Figure 8 displays information plane dynamics using KDE estimation. The GIB formulation (pink) again shows compression phases across all activation functions. Standard IB (blue) fails to show any compression.
F.1.2 Simple Mathematical Functions
Figure 9: Information plane dynamics for simple mathematical functions using KDE estimation. The method provides sharp phase transitions despite increased trajectory variance.
Figure 9 shows KDE-based MI estimation for simple function learning. The GIB formulation exhibits clear compression phases for all functions. Although, this can also be accompanied by a stage of decompression. Standard IB shows more erratic behavior, generally moves upward and leftward but in less distinct phases.
F.1.3 ResNet Information Dynamics
@@
Figure 10: Information plane dynamics for ResNets on CIFAR-10 using KDE estimation.
Figure 10 reveals the first significant divergence in interpretation between the GIB results obtained using binning versus KDE. Under KDE estimation, neither GIB nor IB exhibits interpretable information bottleneck dynamics.
F.1.4 BERT Fine-tuning Dynamics
@@
Figure 11: Information plane dynamics for BERT fine-tuning using KDE estimation.
In Figure 11, our results realign with the interpretation presented in the main text. Following the unlearning phase, we observe a rapid initial fitting step succeeded by a prolonged, gradual (and less pronounced) compression phase. These dynamics are evident for both IB and GIB.
F.2 Gaussian Copula Mutual Information (GCMI)
GCMI (Ince et al., 2017) estimates MI by first transforming variables to have standard Gaussian marginals using the Gaussian copula, then computing MI under the Gaussian assumption. This approach is particularly effective for continuous variables with complex, potentially non-linear relationships. The method is computationally efficient (O(n log n) for n samples) and provides robust estimates even for high-dimensional data. Unlike histogram binning, GCMI does not require discretization parameters and automatically adapts to the data distribution.
F.2.1 Activation Function Comparison
Figure 12: Information plane dynamics across activation functions using GCMI estimation. Standard IB (blue) and GIB (pink) trajectories for networks trained on synthetic tasks S0-S4.
Figure 12 presents information plane dynamics using GCMI estimation across five activation functions. Under this estimator, both methods frequently exhibit compression phases. However, GIB often displays a subsequent decompression phase, characterized by a rightward shift in the later stages of training.
F.2.2 Simple Mathematical Functions
Figure 13: Information plane dynamics for simple mathematical functions using GCMI estimation.
Figure 13 presents results for networks learning arithmetic and polynomial functions using GCMI estimation. While the dynamics of both methods are not easily interpretable, the GIB formulation demonstrates upward leftward movement, indicating the occurrence of fitting and compression but they are occurring at once rather than in distinct phases. In contrast, standard IB shows variable behavior with limited evidence of compression.
F.2.3 ResNet Information Dynamics
Figure 14: Information plane dynamics for ResNets on CIFAR-10 using GCMI estimation.
In Figure 14, we observe that neither the IB nor the GIB yields interpretable results.
F.2.4 BERT Fine-tuning Dynamics
Figure 15: Information plane dynamics for BERT fine-tuning using GCMI estimation.
Figure 15 likewise shows that the dynamics produced by both the IB and GIB are not interpretable.
F.3 Summary of GCMI and KDE Results
In this section, KDE provides strong validation for our main findings. For the KDE method, our GIB formulation consistently shows compression phases where standard IB fails, particularly for different activation functions and simple arithmetic. Meanwhile, the effect is less pronounced for GCMI, but compression remains more likely than for the standard IB. This consistency across three fundamentally different MI estimation approaches (binning, GCMI, and KDE) strongly supports our theoretical framework. The fact that synergy-based decomposition reveals consistent information dynamics across estimation methods suggests that GIB captures a fundamental aspect of how neural networks process information during learning.
@@ Appendix G Computational Complexity
The computational requirements of GIB and IB differ significantly in their scaling behavior. If we define our unit of computation as a single MI estimation, GIB requires $2N+1$ calculations: more specifically, one for the prediction term $I(Z;Y)$ and $2N$ for the complexity term (computing $I(\mathcal{X}^{-i};Q)$ and $I(X^{i};Q)$ for each feature). Critically, these calculations occur at the input layer where dimensionality is typically highest, for CIFAR-10, this means 3072 features. However, GIB’s computational cost is independent of network depth, since it only considers input-output relationships. In contrast, standard IB requires $2L$ MI calculations for $L$ layers, computing $I(\mathcal{X};\mathcal{T}_{l})$ and $I(\mathcal{T}_{l};Y)$ at each layer. While one might compute IB only for the final layer where dynamics are often most pronounced, this prevents the use of the IB as a tool for understanding learning dynamics throughout the network. Additionally, GIB benefits from a key advantage: we can apply PCA to high-dimensional inputs (as we do for CIFAR-10 in Section 5.2) because features at the input layer share a common representation space (Turk and Pentland, 1991). Conversely, combining representations across layers for IB is less conventional.
@@ Appendix H Effect of PCA Dimensionality on GIB Dynamics
Figure 16 demonstrates the impact of PCA dimensionality on observed GIB dynamics. With only 25 principal components, the dynamics are erratic and noisy. However, as we increase to 50 components (shown in main results) and then to 100 components, the dynamics become less noisy. This progression suggests that capturing synergistic information requires sufficient dimensionality to represent the complex feature interactions present in the original input space.